OpenStudy (anonymous):

What is x in 2sinxcosx+2sinx-cosx-1=0

OpenStudy (lastdaywork):

2sinxcosx+2sinx-cosx-1=0 \[2\sin(x)\cos(x)+2\sin(x)-\cos(x)-(\sin^2(x)+\cos^2(x))=0\] Now divide by cos^2(x) on both sides of the equation. Simplify to get a quadratic equation in tan(x). I think the rest is obvious. Note that by dividing with cos^2(x) we implicitly assumed that cos^2(x) ≠ 0 So we need to solve 2sinxcosx+2sinx-cosx-1=0 separately for cos^2(x) = 0

OpenStudy (lastdaywork):

Aah..It can be easily factorized..that makes life a lot easier ;)

OpenStudy (lastdaywork):

@surjithayer I would like to mention that Open Study follows the policy of "Give Help, Not Answers" We should give room to the asker to participate in the process of finding solution. Providing complete solution at one go should be avoided.